OpenStudy (quickstudent):

How do I solve this? 3x + 1 = (1/3)x + 9

OpenStudy (quickstudent):

I'm having trouble finding the first step.

OpenStudy (faiqraees):

Try to bring the variables on one side and constants on the other side

OpenStudy (quickstudent):

How do I do that? Where do I start?

pooja195 (pooja195):

Basically combine like terms Can you identify like terms ?

OpenStudy (quickstudent):

I can identify them. How should I move the like terms when they are on different sides of the = sign? That's what I'm a little confused about

pooja195 (pooja195):

When we want to cancel things out or get them on one side we do the opposite operation. 3x + 1 = (1/3)x + 9 All numbers here are posititve so we will be subtracting I'll get you started off :) We want to isolate x so we will start by putting the x's with the x's and constants with constants Let's start by subtracting 1 from both sides \[\huge~\rm~\bf~3x + 1-1 = (1/3)x + 9-1\] What equation do we have once we subtract 1 from both sides?

OpenStudy (quickstudent):

3x = (1/3)x + 8

pooja195 (pooja195):

Good any ideas on what to do next?

OpenStudy (quickstudent):

Should I subtract 3x from both sides?

pooja195 (pooja195):

Not quite...remember that we want constants on one side and variables on another side

OpenStudy (quickstudent):

How about 1/3x?

pooja195 (pooja195):

What would we do with the 1/3?

OpenStudy (quickstudent):

3x - (1/3)x = (1/3)x + 8 - (1/3)x (2 2/3)x = 8 Does this look alright?

pooja195 (pooja195):

Yes however let's leave it as an improper fraction it'll make things alot more easier ;) So we have \[\huge~\rm~\bf~ \frac{ 8 }{ 3 }x=8\] Any ideas on how we can isolate the x?

OpenStudy (quickstudent):

OK. Hmmm. This? (8/3)x/(8/3) = 8/(8/3) x = 3

pooja195 (pooja195):

Perfect! :)

OpenStudy (quickstudent):

Now I understand! Thanks!:)

pooja195 (pooja195):

You're welcome ^.~